In exercises 43 and 44, use the data set and the indicated number of classes to construct (a) an expanded frequency distribution, (b) a frequency histogram, (c) a frequency polygon, (d) a relative frequency histogram, and (e) an ogive.

43. Pulse Rates

Number of classes: 6

Data set: Pulse rates all students in a class

68 105 95 80 90 100 75 70 84 98 102 70

65 88 90 75 78 94 110 120 95 80 76 108

44. Hospitals

Number of classes: 8

Data set: Number of hospitals in each state

12 100 52 73 354 52 34 8 212 116

13 40 17 142 99 61 76 114 81 50

22 109 56 88 72 16 103 11 28 14

75 37 28 203 156 103 36 176 12 65

27 116 377 35 89 7 62 75 39 13
Answer #1

For data set of Pulse rates,

Number of Observations = n = 24

Number of classes is 6.

Minimum observation is 65 and maximum is 110

110-65=45 , now 45/6= 7.5

Class width should be 8.

So we make Classes

65-73

74-82

83-91

92-100

101-109

110-118

For the data of number of Hospitals,

n=50

Number of classes = 8

Max= 377 , min = 7

Class width= (377-7)/8= 370/8= 46.25

We take 47

Classes are

7-54

55-102

103-150

151-198

199-246

247-294

295-342

343-390